If 22.8% of the unemployed find jobs and 1.2% of the employed lose their jobs, the natural rate of unemployment will be ________ percent (round to the nearest tenth).
Answer #1

Solution:-

Job separation rate (s) = 1.2% = 0.012

Job finding rate (f) = 22.8% = 0.228

Natural rate of unemployment = s / (s + f)

                                                = 0.012 / (0.012 + 0.228)

                                                = 0.012/ 0.24

                                                = 0.05

                                                = 5%